# Norbert Tihanyi

### v1

@@ 2 Motivation
Even when GPT-3.5 is requested to write a secure version of this code –without specifying the vulnerability– it only attempts to verify against entering non-integer inputs by adding the following code snippet: if (scanf("%d", &num1) != 1) {...}. Clearly, after sanitizing the input, the issue of integer overflow is still present. When prompted to create a C program that adds two numbers, it appears that both GPT-3.5 and GPT-4 generate code with this insecure pattern. When asked for a secure version, both models perform input sanitization. By using the ESBMC module to verify this program, the vulnerability is immediately found through a counterexample, and the following message is created:
@@ 3.1 ChatGPT in Software Engineering
Big-Vul
Real-World 188,636 100% ✘/Function CVE/CVW 30 PATCH
Draper
Synthetic+Real-World 1,274,366 5.62% ✘/Function CWE 29 STAT
SARD
Synthetic+Real-World 100,883 100% ✔/Program CWE 114 BDV/STAT/MAN
Juliet
Synthetic 106,075 100% ✔/Program CWE 125 BDV
Devign
Real-World 27,544 46.05% ✘/Function CVE 112 ML
REVEAL
Real-World 22,734 9.85% ✘/Function CVE 32 PATCH
DiverseVul
Real-World 379,241 7.02% ✘/Function CWE 37 PATCH
@@ 5.1 Code generation
To generate the dataset of small C programs, we utilized the GPT-3.5-turbo model [OpenAI(2022)]. We employ GPT-3.5 to generate C code instead of GPT-4 as the latter can be up to $60$ times more expensive than the former model. During the creation process, special attention was given to ensuring the diversity of the FormAI dataset, which contains $112,000$ compilable C samples. Requesting the model to generate a unique C program frequently yields similar outcomes, such as adding two numbers or performing basic string manipulation, which does not align with our objectives. Our focus lies in systematically generating a comprehensive and varied collection of small programs that emulates the code creation process undertaken by developers. Therefore, we need a methodology to circumvent the generation of elementary C programs. To address this issue, we have developed a prompting method consisting of two distinct parts: a dynamic part and a static part. The static component remains consistent and unchanged, while the dynamic portion of the prompt undergoes continuous variation. An example of how a single prompt looks is shown under Figure 4.
Figure 4: Dynamic code generation prompt.
@@ 6.1 CWE classification
23,312 $\mathcal{AO}$
CWE-190, CWE-191, CWE-754, CWE-680, CWE-681, CWE-682
11,088 $\mathcal{ABV}$
CWE-119, CWE-125, CWE-129, CWE-131, CWE-193, CWE-787, CWE-788
88,049 $\mathcal{BO}$
CWE-20, CWE-120, CWE-121, CWE-125, CWE-129, CWE-131, CWE-676, CWE-628, CWE-754, CWE-788
31,829 $\mathcal{DFN}$
CWE-391, CWE-476
24,702 $\mathcal{DFA}$
CWE-119, CWE-125, CWE-125, CWE-131, CWE-129, CWE-755, CWE-787
9823 $\mathcal{DFI}$
CWE-416, CWE-476, CWE-690, CWE-822, CWE-824, CWE-825
5810 $\mathcal{DFF}$
CWE-401, CWE-404, CWE-459, CWE-775
1567 $\mathcal{DZ}$
CWE-369
1620 $\mathcal{O}$
CWE-119, CWE-125, CWE-158, CWE-362, CWE-389, CWE-401, CWE-415, CWE-459, CWE-416, CWE-469, CWE-590, CWE-617, CWE-664, CWE-662, CWE-685, CWE-704, CWE-761, CWE-787, CWE-823, CWE-825, CWE-843

### v3

@@ Abstract
This paper has been published at PROMISE 2023: Proceedings of the 19th International Conference on Predictive Models and Data Analytics in Software. https://doi.org/10.1145/3617555.3617874
@@ 2 Motivation
Even when GPT-3.5 is requested to write a secure version of this code –without specifying the vulnerability– it only attempts to verify against entering non-integer inputs by adding the following code snippet: if (scanf("%d", &num1) != 1) {...}. Clearly, after sanitizing the input, the issue of integer overflow is still present. When prompted to create a C program that adds two numbers, it appears that both GPT-3.5 and GPT-4 generate code with this insecure pattern. When asked for a secure version, both models perform input sanitization. By using the ESBMC module to verify this program, the vulnerability is immediately found through a counterexample, and the following message is created as shown in Figure 3.
@@ 3.1 ChatGPT in Software Engineering
Big-Vul Real-World 188,636 100% ✘/Function CVE/CVW 30 PATCH
Draper Synthetic+Real-World 1,274,366 5.62% ✘/Function CWE 29 STAT
SARD Synthetic+Real-World 100,883 100% ✔/Program CWE 114 BDV/STAT/MAN
Juliet Synthetic 106,075 100% ✔/Program CWE 125 BDV
Devign Real-World 27,544 46.05% ✘/Function CVE 112 ML
REVEAL Real-World 22,734 9.85% ✘/Function CVE 32 PATCH
DiverseVul Real-World 379,241 7.02% ✘/Function CWE 37 PATCH
@@ 5.1 Code generation
To generate the dataset of small C programs, we utilized the GPT-3.5-turbo model [OpenAI(2022)]. We employ GPT-3.5 to generate C code instead of GPT-4 as the latter can be up to $60$ times more expensive than the former model. During the creation process, special attention was given to ensuring the diversity of the FormAI dataset, which contains $112,000$ compilable C samples. Requesting the model to generate a unique C program frequently yields similar outcomes, such as adding two numbers or performing basic string manipulation, which does not align with our objectives. Our focus lies in systematically generating a comprehensive and varied collection of small programs that emulates the code creation process undertaken by developers. Therefore, we need a methodology to circumvent the generation of elementary C programs. To address this issue, we have developed a prompting method consisting of two distinct parts: a dynamic part and a static part. The static component remains consistent and unchanged, while the dynamic portion of the prompt undergoes continuous variation. An example of how a single prompt looks is shown under Listing LABEL:fig:prompt.
Figure 4: Dynamic code generation prompt
@@ 6.1 CWE classification
23,312 $\mathcal{AO}$ CWE-190, CWE-191, CWE-754, CWE-680, CWE-681, CWE-682
11,088 $\mathcal{ABV}$ CWE-119, CWE-125, CWE-129, CWE-131, CWE-193, CWE-787, CWE-788
88,049 $\mathcal{BO}$ CWE-20, CWE-120, CWE-121, CWE-125, CWE-129, CWE-131, CWE-676, CWE-628, CWE-754, CWE-788
31,829 $\mathcal{DFN}$ CWE-391, CWE-476
24,702 $\mathcal{DFA}$ CWE-119, CWE-125, CWE-125, CWE-131, CWE-129, CWE-755, CWE-787
9823 $\mathcal{DFI}$ CWE-416, CWE-476, CWE-690, CWE-822, CWE-824, CWE-825
5810 $\mathcal{DFF}$ CWE-401, CWE-404, CWE-459, CWE-775
1567 $\mathcal{DZ}$ CWE-369
1620 $\mathcal{O}$ CWE-119, CWE-125, CWE-158, CWE-362, CWE-389, CWE-401, CWE-415, CWE-459, CWE-416, CWE-469, CWE-590, CWE-617, CWE-664, CWE-662, CWE-685, CWE-704, CWE-761, CWE-787, CWE-823, CWE-825, CWE-843